OpenStudy (anonymous):

what is the solution to y=3x+1 and y=1/2x- 4 in Coordinate form?

OpenStudy (savvy):

(-2,-5)

OpenStudy (anonymous):

Try 3x+1=1/2 x-4 which will imply 5/2 x = -5 and therefore x = -2

OpenStudy (anonymous):

thnk yooh

OpenStudy (anonymous):

-6x+2y=10 and y=3x+5?

OpenStudy (anonymous):

i jus cnt seem to get it

OpenStudy (savvy):

no one will ever get them...they are parallel lines....

OpenStudy (anonymous):

well i need a method of gettin the anssers iv got a big exam tomorow

OpenStudy (savvy):

ya....just show them that for the 2 lines the ratio of the coeff. of x of both the lines is equal to the ratio of the coeff. of y for both of them.....hence the lines have to be parallel.....

OpenStudy (anonymous):

ok now whats the solution?

OpenStudy (savvy):

the solution is that - "the lines are parallel and hence will never meet so there is no solution to the equations"

OpenStudy (anonymous):

o thnk yooh lol

OpenStudy (savvy):

you may remember the following: for 2 linear equations a1x + b1y + c1 = 0 a2x + b2y + c2 = 0 if, (a1/a2) is not equal to (b1/b2), then there exists a single/unique solution. if, (a1/a2) = (b1/b2) not equal to (c1/c2), then lines are parallel and no solution exists. if, (a1/a2)=(b1/b2)=(c1/c2), then lines are coincident and infinite solutions exist.